1) Five times the difference of a number and two is seven more than that number
2) Twice a number decreased by four is greater than ten, what numbers satisfy this condition?
3) Twice a number decreased by two is equivalent to than number increased by five.

Answer by Edwin McCravy(20077)

1) Five times the difference of a number and two is seven more than that number
5(N-2) = N+7

You solve!


2) Twice a number decreased by four is greater than ten, what numbers satisfy this condition?
2N-4 > 10

  2N > 16

   N > 8

All numbers greater than 8 satisfy this condition.

3) Twice a number decreased by two is equivalent to than number increased by five.
2N-2 = N+5

You solve.

Edwin